OpenStudy (buttermequeasy):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (nottim):

All of them at the same time?

OpenStudy (anonymous):

so perpendicular means has the opposite resiprical slope,do u know how to find that?

OpenStudy (nottim):

That perpendicular.

OpenStudy (nottim):

Not accurate drawing by the way.

OpenStudy (lgbasallote):

point slope form probably is the best here...according to your equation...the slope is -1..however you want perpendicular so engative reciprcocal is 1 plug it in the point slope equation y - y1 = m(x-x1) so y - (-2) = 1(x- (-4)) so y+ 2 = x+4 equation of perpendicular line is y = x + 2 in slope-intercept form in general form...it's x - y +2 doesnt make any difference